OpenStudy (anonymous):

simplify x+3/9 / x+12/6

OpenStudy (anonymous):

(x + 3)/9 / (x + 12)/6 = (x + 3)6/(x+12)9

OpenStudy (kinggeorge):

\[\Large {{x+3 \over 9} \over {x+12 \over 6}} ={6(x+3) \over 9(x+12)}\]

OpenStudy (anonymous):

you can do the rest